# First Heuristic Then Rational: Dynamic Use of Heuristics in Language Model Reasoning

### v1

@@ Abstract
Multi-step reasoning is widely adopted in the community to explore the better performance of language models (LMs). We report on the systematic strategy that LMs use in this process. Our controlled experiments reveal that LMs rely more heavily on heuristics, such as lexical overlap, in the earlier stages of reasoning when more steps are required to reach an answer. Conversely, as LMs progress closer to the final answer, their reliance on heuristics decreases. This suggests that LMs track only a limited number of future steps and dynamically combine heuristic strategies with logical ones in tasks involving multi-step reasoning. [^1] The code/data will be made public upon acceptance.
@@ 1 Introduction
When facing complex tasks, humans tend to seek shallow, heuristic solutions first; and, once these attempts are revealed to fail or elicit another reasonable solution, they switch to being more rational Erickson and Mattson (1981); Frederick (2005). This systematic behavior helps us to predict how humans will tackle new problems. Given such a view, when it comes to predicting the behavior of language models (LMs) Madaan and Yazdanbakhsh (2022); Ye et al. (2023), the following question naturally arises—Do LMs also use a similar systematic strategy to solve complex tasks, or is their strategy totally different from humans, or do they have no such strategies? This study explores an answer to this question. Investigating LMs’ strategic behavior in problem-solving is expected to provide a new perspective on LMs’ reasoning mechanism. It may also address general concerns that current neural models tend to overly rely on superficial, heuristic cues and may end up with irrational conclusions Du et al. (2022); Lai et al. (2021); Jia and Liang (2017); Ye et al. (2023); Chen et al. (2024).
Specifically, in the course of step-by-step reasoning, they tend to rely on shallow, short-sighted heuristic preferences in choosing rules in the early phase and dynamically switch to be more rational and goal-oriented to make the right choice to reach the goal. This highlights a severe limitation of present LMs, including GPT-4 OpenAI (2023), in searching for a solution to multi-hop reasoning tasks, particularly when tasks require many-step long solutions.
Figure 1: Illustration of the systematic strategy we discovered in language models (LMs). When the goal is distant from the current state in a multi-step reasoning process, the models tend to rely on heuristics, such as superficial overlap, which can lead them in the wrong direction. In contrast, when the goal is within a limited distance, the models are more likely to take rational actions to reach the goal.
Figure 2: Overview of the task setting. Given premises and a question, a model answers the question step-by-step (left part). Through each reasoning step $t$ of selecting/paraphrasing relevant premise $p_{k}\in P$, the available facts $\bm{z}$ are enriched, and if that step is necessary to reach the answer, the distance to the answer $d$ decreases (right part).
@@ 2 Task
As a controlled testbed to analyze LM’s reasoning ability, we adopt an arithmetic reasoning task (Figure 2 left). We will use both natural and artificially controlled datasets in the experiments, but let us use the latter, more formal examples to explain the task overview.
@@
The problem consists of a set of premises $P=\{p_{1},\cdots,p_{k}\}$ and a question $q$. Each premise describes either type of fact: (i) Person A has $n$ items (A=$n$), or (ii) Person B has $n$ more/less items than A has (B=A+$n$ or B=A-$n$). The question asks the exact number of certain items a particular person ultimately has (How many items does B have?). Here, one should consider multiple premises to derive the final answer, e.g., A=3;B=2+A;B=2+3=5. Notably, some premises are irrelevant to the answer; thus, models have to track which premise is necessary to reach the final answer.
@@
Let $f$ be a model that is instructed to solve the task step-by-step. In each reasoning step $t$, the model $f$ selects a particular premise $p_{i}\in P$ and paraphrases it into a new fact $z_{t}$ by eagerly resolving reference expressions based on the already stated facts $\bm{z}_{<t}=[z_{1},\cdots,z_{t-1}]$:
$f(P,q,\bm{z}_{<t})=(p_{i},z_{t})\;\;\,\mathrm{.}$ (1)
For example, in Figure 2, when $p_{2}$ “Walter has 2 more apples than Peggy.” is selected at a particular reasoning step, the respective $z_{t}$ should be “Walter has 2+5=7 apples.” if $\bm{z}_{<t}$ already contains the number of apples Peggy has, i.e., $p_{1}$. [^2] If the reference can not be resolved with $\bm{z}_{<t}$, the model just repeats the selected premise $p_{i}$ as $z_{t}$. Starting with an empty set of stated facts $\bm{z}=\{\}$, the model recursively performs a reasoning step and can stop when outputting a special symbol EOS or providing an answer to the question $q$. Here, we denote the whole history of selected premises as $\bm{h}=[p_{i},\cdots,p_{j}]\in P^{*}$, where $P^{*}$ is Kleene closure of $P$. Its $t$-th element $h_{t}$ is the premise employed to derive the $t$-th reasoning step. Henceforth, we call $\bm{h}$ reasoning steps and focus on the ability to search the right $\bm{h}$.
@@
Each problem has a set of solutions $H^{\circ}\subset P^{*}$. Specifically, the final stated fact $z_{-1}$ generated from the final reasoning step $h_{-1}$ of the solution $\bm{h}\in H^{\circ}$ should provide a concrete number asked by the question $q$. Figure 2 illustrates such a set of solutions $H^{\circ}$ as the steps leading to the final states of the state transition graph (right part of Figure 2).
@@
Within the set of solutions, there is only one minimal solution $\bm{h}^{*}\in H^{\circ}\subset P^{*}$. To define $\bm{h}^{*}$, let us first introduce a distance to the answer. In each reasoning step $t$, one can determine the minimum number of remaining reasoning steps to reach the answer $d\in\mathbb{N}$, given $\bm{h}_{\leq t}\in P^{*}$ and the initially provided premises $P$. Intuitively, $d$ can be derived from a state transition graph and the minimum number of transitions to the closest final states, as shown in Figure 2 (right part). Here, we denote the mapping function from $\bm{h}_{\leq t}$ to $d$ as $g:P^{*}\rightarrow\mathbb{N}$. For example, $g([p_{2},p_{1},p_{2}])=1$ in Figure 2. Minimal solution $\bm{h}^{*}$ satisfies $\forall t\ g(\bm{h}^{*}_{\leq t})<g(\bm{h}^{*}_{\leq t-1})$; that is, $\bm{h}^{*}$ does not contain any irrelevant step to approach the answer. For example, the MS reasoning steps in Figure 2 are $[p_{1},p_{2},p_{4}]=\bm{h}^{*}$.
@@
We evaluate LMs’ ability to derive the minimal solution $\bm{h}^{*}$ as instructed by 4-shot examples 10. Notably, we do not care about the ability to correctly introduce a new fact $z_{t}$ (Eq. 1), e.g., the accuracy of arithmetic operation (e.g., 5+2=7), but separately focus on their search strategy to select the relevant premise to perform the next reasoning step.
@@ 3 Heuristics
Whether models tend to select premises with the same person name (PN) as the one in question is targeted. For example, given a question “how many apples Judy has,” premises such as “Judy’s mother got 3 apples” can be a distractor when it is unnecessary to reach the answer.
@@
Whether models tend to select the premise in the initial position of context is targeted.
Grammatical feature of premise (Neg.):
Given that a specific grammatical feature, e.g., whether the sentence has a negation word, is often a superficial cue Du et al. (2021); Niven and Kao (2019), we specifically analyze the bias that models avoid selecting premise with negation.
We append distractor premises that are irrelevant to answering the question but presumably preferred by the model w.r.t. some heuristics to the input $P$. We analyze whether LMs select such distractors to understand their search strategy (§ 4).
@@ 4 Experiments
We use four representative variants of LLMs: text-bison-001 version of Google’sPaLM2 Anil et al. (2023), Llama2-13B Touvron et al. (2023), gpt-3.5-turbo-0125 and gpt-4-0613 snapshots of OpenAI’s GPT-3.5-turbo OpenAI (2022) and GPT-4 OpenAI (2023).
@@ 4.1 Preliminary experiments
First, we confirm that LLMs exploit specific heuristics during step-by-step reasoning in natural and artificially-controlled datasets.
@@
We use two datasets: GSM8K Cobbe et al. (2021) (App. A) and artificially-controlled data with 4-step arithmetic reasoning (App. B). Each instance in either dataset has at least one premise $\tilde{p}\not\in\bm{h}^{*}$ that is irrelevant to inducing the answer (i.e., distractor).
We randomly select one such irrelevant premise (Base), [^3] Strictly speaking, we randomly selected the premise that does not match any of the three heuristics. modify it to match a particular heuristics (e.g., move it to the first position), and observe whether such a modification makes it more attractive for LLMs to select during step-by-step reasoning. We report the percentage of the problems, where the model selected the distractor at least once during reasoning. We separately run the experiments for each type of heuristic rather than adding all types of distractors to the input at once.
@@
Table 1: The percentage of the problems where the model selected a distractor $\tilde{p}$ in step-by-step reasoning. Over. and Pos. denote the Overlap and Position biases.
@@ Results:
Table 1 shows the results. The Base results (baseline) indicate how many times the randomly sampled distractor $\tilde{p}$ is selected during reasoning. The scores for the distractors with heuristic features (Over., Pos., Neg. in Table 1) are generally higher (lower for Neg.) than the Base scores across models and datasets. This indicates that LMs, on average, tend to rely on our targeted heuristics (§ 3). Note that, interestingly, different models yield somewhat different preference towards distractor types; for example, Llama2 and GPT-3.5 are more biased premise position than PaLM2 and GPT-4.
@@ 4.2 Main experiments
Then, we further investigate the strategies of LMs to employ heuristics. We hypothesized that the more distant the reasoning step is from the answer (higher $d$), the more heavily models rely on heuristics. This is motivated by the larger gap between questions and available knowledge, leading to difficulty. Thus, models have to rely on more primitive and heuristic factors.
@@
Table 2: Example of a distractor examined in § 4.2. Suppose that $h^{*}_{1}$ is “Peggy has 5 apples.” One of the two candidates with “Peggy” is the correct next step $h^{*}_{2}$ (green), and the other is a distractor (orange).
Figure 3: How frequently a particular distractor is selected (y-axis: $r$) in each reasoning step (x-axis: $d$).
@@
To identify in which steps heuristics are more likely to be exploited, ideally, one should design a distractor equally attractive to all the reasoning steps in $\bm{h}^{*}$ and analyze when it is selected during the reasoning to facilitate a fair step-wise comparison; however, such a distractor is inherently difficult to implement. Instead, we add multiple distractors $\tilde{P}$ to the artificial data; each of them $\tilde{p}_{t}\in\tilde{P}$ correspond to each reasoning step $h^{*}_{t}$ in the sense that both share the same person name that appeared in the previous step $h^{*}_{t-1}$ (Table 2). [^4] To rule out the shortcut cue regarding the reference frequencies of each person name, we further added distractor premises to make the frequencies uniform. Similar to § 4.1, we further modify each distractor $\tilde{p}_{t}\in\tilde{P}$ to match each heuristic (Overlap with question or Position in § 3). [^5] We excluded the Neg. (avoidance) bias because if a model avoids negation in the latter step, we could not distinguish whether it was due to heuristic or rational search. In evaluation, for each $t$, partial correct reasoning steps $\bm{h}^{*}_{<t}$ are teacher-forced to a model, and we analyze whether the model selects the right next step $h^{*}_{t}$ or its respective distractor $\tilde{p}_{t}$. We calculate the frequency #()#\#($\cdot$)# ( ⋅ ) of models’ selecting $\tilde{p}_{t}$ and $h^{*}_{t}$; then, their ratio $r=\frac{\#\tilde{p}_{t}}{\#h^{*}_{t}}$ is reported. The chance rate (i.e., random premise selection) is 0.5. Our hypothesis is that the more current step $t$ is distant from the goal (i.e., the larger $d=g(\bm{h}^{*}_{t})$), the more frequently the distractor is selected (i.e., the higher $r$).
@@
We used 5-step artificial reasoning data and excluded the first ($d=5\rightarrow 4$) and the last ($d=1\rightarrow 0$) steps from evaluation regarding their special properties. GSM8K was excluded due to the infeasibility of controlled distractors.

### v2

@@ Abstract
Multi-step reasoning instruction, such as chain-of-thought prompting, is widely adopted to explore better language models (LMs) performance. We report on the systematic strategy that LMs employ in such a multi-step reasoning process. Our controlled experiments reveal that LMs rely more heavily on heuristics, such as lexical overlap, in the earlier stages of reasoning, where more reasoning steps remain to reach a goal. Conversely, their reliance on heuristics decreases as LMs progress closer to the final answer through multiple reasoning steps. This suggests that LMs can backtrack only a limited number of future steps and dynamically combine heuristic strategies with rationale ones in tasks involving multi-step reasoning. [^1] The code/data is available in https://github.com/ao1neko/Heuristic-and-Rational-Reasoning
@@ 1 Introduction
When facing complex tasks, humans tend to seek shallow, heuristic solutions first Erickson and Mattson (1981); Frederick (2005). Once these attempts are revealed to fail or elicit another reasonable solution, they switch to being more rational Stanovich and West (2000). Such systematic behavior helps us to predict how humans will tackle new problems. Given such a view, when it comes to predicting the behavior of language models (LMs) (Madaan and Yazdanbakhsh 2022; Zhang et al. 2024; inter alia), the following question naturally arises—Do LMs also use a similar systematic strategy to solve complex tasks, or is their strategy totally different from humans, or do they have no such strategies? This study explores an answer to this question. Such analyses will shed light on the cognitive plausibility of LMs in problem solving Opedal et al. (2024); Eisape et al. (2024); Aher et al. (2023) as well as address general concerns of current neural models relying on superficial, heuristic cues overly and ending up with irrational conclusions Du et al. (2022); Lai et al. (2021); Jia and Liang (2017); Ye et al. (2023); Chen et al. (2024).
In this paper, we demonstrate that LMs rely on shallow heuristics more frequently in the earlier phase of multi-step reasoning, and then gradually switch their reasoning strategy to be more rational and goal-oriented to make the right choice to reach the goal. From an engineering perspective, this highlights a limitation of modern LMs, including GPT-4 OpenAI (2023), in searching for a solution at the initial stage of step-by-step reasoning, particularly when tasks require many-step-long solutions, implying that they can backtrack only a limited number of future steps from the answer to the current progress of reasoning. From a cognitive perspective, their behaviors would be somewhat human-like in the sense that the models try to employ heuristics first in solving a complex problem. Moreover, this paper is the first to show that models are equipped with both heuristic and goal-oriented reasoning and dynamically switch between them as needed.
Figure 1: Illustration of the systematic reasoning strategy we discovered in language models Refer to caption. When the goal is distant from the current reasoning step, they tend to rely on heuristics Refer to caption to take the next reasoning step, such as lexical overlap with a question, leading to the wrong direction (red path). In contrast, when the goal is within a limited distance, they are more likely to take rational actions (green path) to reach the goal.
Figure 2: Overview of the task setting. Given premises and a question, a model answers the question step-by-step (left part). Through each reasoning step $t$ of selecting/paraphrasing relevant premise $p_{k}\in P$, the available facts $\bm{z}$ are enriched (reasoning state progresses in the right part). If a reasoning step follows the minimal solution (green path in the right part), the distance to the answer $d$ decreases.
@@ 2 Task
We adopt an arithmetic reasoning task as a controlled testbed to analyze LM’s reasoning ability (Figure 2 left). We will use natural and artificially controlled datasets in the experiments, but let us use the latter, more formal examples to explain the task overview.
@@
The problem consists of a set of premises $P=\{p_{1},\cdots,p_{k}\}$ and a question $q$. Each premise describes either type of fact: (i) Person A has $n$ items (A=$n$), or (ii) Person B has $n$ more/fewer items than A has (B=A+$n$ or B=A-$n$). The question $q$ asks the exact number of specific items a particular person ultimately has (How many items does B have?). Here, one should consider multiple premises to derive the final answer, e.g., A=3; B=2+A; B=2+3=5. Notably, some premises are irrelevant to the answer; thus, models have to track which premise is necessary to reach the final answer.
@@
Let $f$ be a model that is instructed to solve the task step-by-step. In each reasoning step $t$, the model $f$ selects a particular premise $p_{i}\in P$ and paraphrases it into a new fact $z_{t}$ by eagerly resolving reference expressions based on the already stated facts $\bm{z}_{<t}=[z_{1},\cdots,z_{t-1}]$ as in equation (1):
$(p_{i},z_{t})=f(P,q,\bm{z}_{<t})\;\;\,\mathrm{.}$ (1)
For example, in Figure 2, when $p_{2}$, Walter has 2 more apples than Peggy., is selected at a particular reasoning step, the respective $z_{t}$ should be Walter has 2+5=7 apples. if $\bm{z}_{<t}$ already contains the number of apples Peggy has, i.e., $p_{1}$. [^2] If the reference can not be resolved with $\bm{z}_{<t}$, the model repeats the selected premise $p_{i}$ as $z_{t}$. Starting with an empty set of stated facts $\bm{z}=\{\}$, the model recursively performs a reasoning step and can stop when outputting a special symbol EOS or answering the question $q$. Here, we denote the whole history of selected premises as $\bm{h}=[p_{i},\cdots,p_{j}]\in P^{*}$, where $P^{*}$ is Kleene closure of $P$. Its $t$-th element $h_{t}$ is the premise to derive the $t$-th reasoning step $z_{t}$. Henceforth, we call $\bm{h}$ reasoning steps and focus on the ability to search for the right $\bm{h}$.
@@
Among the possible reasoning steps $P^{*}$, there is a set of solutions $H^{\circ}\subset P^{*}$, where the final stated fact $z_{-1}$ in a solution $h\in H^{\circ}$ yields the right answer to the question $q$. Figure 2 illustrates such a set of solutions $H^{\circ}$ as the steps leading to the final states of the state transition graph (right part of Figure 2), e.g., $[p_{1},p_{3},p_{2},p_{4}]\in H^{\circ}$.
@@
Within the set of solutions, there is only one minimal solution $\bm{h}^{*}\in H^{\circ}\subset P^{*}$. Intuitively, $\bm{h}^{*}$ does not contain any irrelevant step to approach the answer; for example, the minimal solution of the problem in Figure 2 is $[p_{1},p_{2},p_{4}]=\bm{h}^{*}$. To define $\bm{h}^{*}$, let us first introduce a distance to the answer. In each reasoning step $t$, one can determine the minimum number of remaining reasoning steps to reach the answer $d\in\mathbb{N}$, given $\bm{h}_{\leq t}\in P^{*}$ and the initially provided premises $P$. The distance $d$ can be derived from a state transition graph and the minimum number of transitions to the closest final states, as shown in Figure 2 (right part). Here, we denote the mapping function from $\bm{h}_{\leq t}$ to $d$ as $g:P^{*}\rightarrow\mathbb{N}$. For example, $g([p_{2},p_{1},p_{2}])=1$ in Figure 2. A minimal solution $\bm{h}^{*}$ satisfies $\forall t\ g(\bm{h}^{*}_{\leq t})<g(\bm{h}^{*}_{\leq t-1})$.
@@
We evaluate LMs’ ability to derive the minimal solution $\bm{h}^{*}$ as instructed by 4-shot examples (Table 10). Notably, we do not care about the ability to correctly introduce a new fact $z_{t}$ (Eq. 1), e.g., the accuracy of arithmetic operation (e.g., 5+2=7), but separately focus on their search strategy to select the relevant premise to perform the next reasoning step.
@@ 3 Heuristics
Neural models generally tend to rely on superficial, shallow similarity of texts when considering their associations. We specifically examine whether models select premises with the same person name (PN) as the one in question as a representative of such biases. For example, given a question how many apples Judy has, premises such as Judy’s mother got 3 apples might be selected as a relevant fact, regardless of its necessity to reach the answer.
@@
It has been reported that models tend to select information, e.g., first and last, in specific positions in the context Liu et al. (2024). We examine whether models tend to select the premise in the initial position of context.
Grammatical feature of premise (Negative):
Given that a specific grammatical feature, e.g., negation word, is often a superficial cue Du et al. (2021); Niven and Kao (2019), we specifically analyze the bias that models avoid selecting premise with negation word, i.e., not.
@@ 4 Experiments
We first confirm that models indeed rely on particular types of heuristics in our setting (§ 4.1). Then, we investigate when in the step-by-step reasoning, such heuristics are more frequently exploited (§ 4.2).
General settings:
We use four representative variants of large language models (LLMs): text-bison-001 version of Google’sPaLM2 Anil et al. (2023), Llama2-13B Touvron et al. (2023), gpt-3.5-turbo-0125 and gpt-4-0613 snapshots of OpenAI’s GPT-3.5-turbo OpenAI (2022) and GPT-4 OpenAI (2023). These models are instructed to yield a minimum solution via prompting.
@@ 4.1 Preliminary experiments
First, we confirm that LLMs exploit specific heuristics in natural and artificially controlled datasets during step-by-step reasoning.
@@
We use two datasets as examples of multi-hop reasoning: GSM8K Cobbe et al. (2021) (App. A) and artificially-controlled dataset with 4-step arithmetic reasoning (App. B). To perform controlled experiments towards the LLMs’ use of heuristics (§ 4), we extend GSM8K and the artificially controlled dataset to the Overlap, Possition, and Negative variants. Each variant is created by adding one premise; there, the use of its corresponding heuristics make the model fail to find the minimal solution in step-by-step reasoning (1). For example, the Position dataset has an additional premise $\tilde{p}$ (i.e., distractor) at the beginning of the sentences, which is irrelevant to answering the question $\tilde{p}\not\in\bm{h}^{*}$. The Position heuristics will lead the model to select the first sentence, i.e., distractor, as a neccesarry fact. As a baseline, we also created Base by adding a random distractor that does not match any of the three heuristics. If the model more frequently selects the distractors in the Overlap, Possition, and Negative datasets, compared to Base, we can confirm that models are at least biased towards our selected heuristics. We describe the details of the dataset creation process in Appendices §§ A.1 and B.1.
@@
Table 1: The frequency of the problems where the model selected a distractor $\tilde{p}$ in step-by-step reasoning. “Over.”, “Pos.” and “Neg.” denote the Overlap, Position, and Negative heuristics. The results are bolded when the frequency of output misled by the distractor increased (Over.↑ and Pos.↑) or decreased (Neg.↓) compared to the base results. In camera-ready, add the above explanation to the caption.
@@ Results:
Table 1 shows the frequencies of mentioning the distractor at least once during reasoning. The scores for Overlap, Position, and Negative in Table 1) are generally higher (lower for Negative) than the Base scores across models and datasets. This indicates that LLMs, on average, tend to rely on our targeted heuristics (§ 3). Interestingly, different models yield different preferences towards distractor types; for example, Llama2 and GPT-3.5 have more biased premise positions than PaLM2 and GPT-4. Note that whether or not a distractor was selected was determined by some rules. Such details are exmplained in the Appendices §§ A.2 and B.2.
@@ 4.2 Main experiments
Then, we further investigate the LMs’ dynamic use of heuristics. We hypothesized that the more distant the current reasoning step is from the answer (higher $d$ in § 2), the more heavily models rely on heuristics. Again, generating the minimal solution requires the model to track/plan the future path (remaining necessary and sufficient information) to reach the final answer, and its remaining length becomes longer at the initial phase of reasoning. If models have a limited capacity to track the future path, they may have to give up rational reasoning and rely on heuristics, particularly at the earlier stage of reasoning, where the volume of the remaining future path is likely to exceed the model’s capacity.
@@
Table 2: Example of a distractor examined in § 4.2. Suppose that $h^{*}_{1}$ is “Peggy has 5 apples.” Two candidate premises with “Peggy” seem to be plausible continuations as $h^{*}_{2}$, but only one is relevant to the final answer (green), and the other is a distractor (orange).
Figure 3: The ratio at which a particular distractor is selected (y-axis: $r$) in each reasoning step (x-axis: $d$).
@@
To investigate the relationship between the distance to the answer and the models’ reliance on heuristics, we identify at which steps heuristics are more likely to be exploited. Ideally, to facilitate a fair step-wise comparison, one should design a distractor equally attractive to all the reasoning steps in $\bm{h}^{*}$ and analyze when it is selected during the reasoning; however, such a distractor is inherently difficult to implement. Instead, we prepare multiple distractors $\tilde{P}$ to the problem in artificial data; each of them $\tilde{p}_{t}\in\tilde{P}$ correspond to each reasoning step $h^{*}_{t}$ in the sense that both share the same person name that appeared in the previous step $h^{*}_{t-1}$ (Table 2). [^3] To rule out the shortcut cue regarding the reference frequencies of each person name, we further added distractor premises to make the frequencies uniform. Similar to § 4.1, we further modify each distractor $\tilde{p}_{t}\in\tilde{P}$ to match each heuristic (Overlap with question or Position in § 3). [^4] We excluded the Negative (avoidance) bias because if a model avoids negation in the latter step, we cannot distinguish whether it was due to the heuristic or rational search. In evaluation, for each $t$, partial correct reasoning steps $\bm{h}^{*}_{<t}$ are teacher-forced to a model, and we analyze whether the model selects the right next step $h^{*}_{t}$ or its respective distractor $\tilde{p}_{t}$. Specifically, we calculate the frequency #()#\#($\cdot$)# ( ⋅ ) of models’ selecting $\tilde{p}_{t}$ or $h^{*}_{t}$; then, the ratio of exploiting a distractor $r=\frac{\#\tilde{p}_{t}}{\#\tilde{p}_{t}+\#h^{*}_{t}}$ is reported. The chance rate should be 0.5.
Note that we could not use the GSM8K dataset in this experiment since designing such controlled distractors for each reasoning step was not feasible, i.e., our creation of artificial, controlled data enables this kind of analysis. In addition, we will not analyze the Negative heuristic in this experiment because it is a bias in the direction of avoidance, making the experimental design complicated.
@@
We used 5-step artificial reasoning data; that is, the distance to the answer is five at first $d=5$ and will monotonically decrease ($d=5\rightarrow 4\rightarrow\cdots\rightarrow 0$) through the steps in the minimal solution. The first ($d=5\rightarrow 4$) and the last ($d=1\rightarrow 0$) steps are excluded from our analysis for their special properties; e.g., the right last step can be identified simply by detecting the lexical overlap with the question.